a, b, c are 3 vectors. a×b=b×c not equal to zero.
a+c=?

Grade:12

1 Answers

Khimraj
6 years ago
a, b and c are vector
and axb = bxc
 - axb + 0 + bxc =0
bxa + bxb + bxc =0
bx(a+b+c) = bx0
So a+b+c =0
then a+c = -b.
Hope it clears. If you like answer then please approve it.